# Generating coupled cluster code for modern distributed memory tensor software

### v1

@@ IV.3 Application: CO
The study of the CO molecule was chosen as an application, due to the interesting aspects of its triple bond stretching, studied in a recent work by the Shabaev group.Usov et al. (2024) There, the influence of quantum electrodynamics (QED) corrections is analyzed which is relevant for the HAMP-vQED projectSaue et al. (2024) for which the present code was created. These are small-scale effects and their study requires highly accurate methods.
We used the recently implemented amfX2C (XAMFI) correctionKnecht et al. (2022) for relativistic picture-change effects.
@@
As expected, single-reference methods struggle for large bond distances. The nonparallelity of CCSD and CCSDT becomes apparent at the bond distance of Å (Fig. 7) and the CC has convergence issues from Å on. This is also visible for the rightmost two points for CCSDTQ in Fig. 8.
A deeper study of the behavior of higher-order CC upon dissociation, and especially finding the minimum CC excitation level to break the bond successfully is beyond the scope of this manuscript, and will be a subject for our further work. Furthermore, we will study how amfX2C and QED corrections influence the result.
@@ V CONCLUSIONS
The code generated from tenpi represents the first implementation of CCSDT and CCSDTQ in DIRAC. For massively parallel CCSD, it shows a performance comparable to the hand-tuned code with the measured overhead of only 30%, which is an improvement over the 80% overhead of the existing generated CCSD.
Overall, tenpi provides users with an elegant and systematic way of implementing the coupled cluster methods while hiding under the hood most of the complexity related to the equations, parallelization and intermediates. A simple python interface is featured, while the output is a highly optimized Fortran code that controls the parallel tensor library.
@@
The outlook for tenpi includes the development of CCSD(T) and CCSDT(Q) for distributed-GPU platforms, implementation of spatial-symmetry support, and the integration of DIRAC with new tensor software. Improvement of intermediate optimization for these applications can studied and the impact on internode communication minimized. The method development can be boosted by a symbolic interface to Mathematica. tenpi can also serve as a standalone production CC package, if the integral generation is interfaced directly to an existing code, like ReSpect.Repisky et al. (2020)
@@ Acknowledgements.
- Dongarra (2023) J. Dongarra, “A Not So Simple Matter of Software,” (2023).
@@
- Woolston (2022) C. Woolston, “Why science needs more research software engineers,” (2022).
- Lyakh (2019) D. I. Lyakh, International Journal of Quantum Chemistry 119, (2019).
@@
- Papadopoulos (2019) A. Papadopoulos, “Necessity of Code Generators in Electronic-Structure Theories and their Infrastructure,” (2019).
@@
- The Mathworks, Inc. (2024) The Mathworks, Inc., “MATLAB,” (2024).
- Waterloo Maple, Inc. (2024) Waterloo Maple, Inc., “Maple,” (2024).
- Inc. (2024) W. R. Inc., “Mathematica, Version 14.0,” (2024), champaign, IL, 2024.
- The Khronos Group, Inc. (2021) The Khronos Group, Inc., “SYCL,” (2021).
- The OpenACC Organization (2022) The OpenACC Organization, “OpenACC,” (2022).
@@
- Brandejs et al. (2024) J. Brandejs, P. Bientinesi, L. Visscher, A. Gomes, T. Saue, et al., “CECAM Workshop on Tensor Contraction Library Standardization,” (2024).
@@
- Shiozaki (2017) T. Shiozaki, WIREs Computational Molecular Science 8, (2017).
@@
- Song, Henderson, and Scuseria (2022) R. Song, T. M. Henderson, and G. E. Scuseria, The Journal of Chemical Physics 156, (2022).
- Saitow, Kurashige, and Yanai (2013) M. Saitow, Y. Kurashige, and T. Yanai, The Journal of Chemical Physics 139, (2013).
- Neuscamman, Yanai, and Chan (2009) E. Neuscamman, T. Yanai, and G. K.-L. Chan, The Journal of Chemical Physics 130, (2009).
- Köhn, Richings, and Tew (2008) A. Köhn, G. W. Richings, and D. P. Tew, The Journal of Chemical Physics 129, (2008).
@@
- Quintero-Monsebaiz and Loos (2023) R. Quintero-Monsebaiz and P.-F. Loos, AIP Advances 13, (2023).
- Lyakh, Ivanov, and Adamowicz (2004) D. I. Lyakh, V. V. Ivanov, and L. Adamowicz, The Journal of Chemical Physics 122, (2004).
@@
- Yadav et al. (2024) R. Yadav, M. Bauer, D. Broman, M. Garland, A. Aiken, and F. Kjolstad, “Automatic Tracing in Task-Based Runtime Systems,” (2024).
@@
- Psarras et al. (2021) C. Psarras, L. Karlsson, J. Li, and P. Bientinesi, “The landscape of software for tensor computations,” (2021).
@@
- Lyakh et al. (2022) D. I. Lyakh, T. Nguyen, D. Claudino, E. Dumitrescu, and A. J. McCaskey, Frontiers in Applied Mathematics and Statistics 8, (2022).
- Sharir, Shashua, and Carleo (2022) O. Sharir, A. Shashua, and G. Carleo, Physical Review B 106, (2022).
@@
- Springer and Bientinesi (2016) P. Springer and P. Bientinesi, “Design of a high-performance GEMM-like Tensor-Tensor Multiplication,” (2016).
- cuT (2019) “NVIDIA cuTENSOR: A High-Performance CUDA Library For Tensor Primitives,” (2019).
- Bientinesi et al. (2022) P. Bientinesi, D. Ham, F. Huang, P. H. J. Kelly, P. S. Sadayappan, and E. Stow, Dagstuhl Reports , (2022).
@@
- Murray et al. (2023) R. Murray, J. Demmel, M. W. Mahoney, N. B. Erichson, M. Melnichenko, O. A. Malik, L. Grigori, P. Luszczek, M. Dereziński, M. E. Lopes, T. Liang, H. Luo, and J. Dongarra, “Randomized Numerical Linear Algebra : A Perspective on the Field With an Eye to Software,” (2023), arXiv.
@@
- tal (2018) “TAL-SH,” (2018).
- Mutlu et al. (2023) E. Mutlu, A. Panyala, N. Gawande, A. Bagusetty, J. Glabe, J. Kim, K. Kowalski, N. P. Bauman, B. Peng, H. Pathak, J. Brabec, and S. Krishnamoorthy, The Journal of Chemical Physics 159, (2023).
@@
- Visscher et al. (2024) L. Visscher, H. J. Aa. Jensen, R. Bast, A. S. P. Gomes, T. Saue, I. A. Aucar, V. Bakken, J. Brandejs, C. Chibueze, J. Creutzberg, K. G. Dyall, S. Dubillard, U. Ekström, E. Eliav, T. Enevoldsen, E. Faßhauer, T. Fleig, O. Fossgaard, L. Halbert, E. D. Hedegård, T. Helgaker, B. Helmich-Paris, J. Henriksson, M. van Horn, M. Iliaš, C. R. Jacob, S. Knecht, S. Komorovský, O. Kullie, J. K. Lærdahl, C. V. Larsen, Y. S. Lee, N. H. List, H. S. Nataraj, M. K. Nayak, P. Norman, A. Nyvang, M. Olejniczak, J. Olsen, J. M. H. Olsen., A. Papadopoulos, Y. C. Park, J. K. Pedersen, M. Pernpointner, J. V. Pototschnig, R. Di Remigio Eikås, M. Repiský, K. Ruud, P. Salek, B. Schimmelpfennig, B. Senjean, A. Shee, J. Sikkema, A. Sunaga, J. Thyssen, J. van Stralen, M. L. Vidal, S. Villaume, O. Visser, T. Winther, S. Yamamoto, and X. Yuan, “DIRAC24,” (2024).
- Saue et al. (2020) T. Saue, R. Bast, A. S. P. Gomes, H. J. Aa. Jensen, L. Visscher, I. A. Aucar, R. Di Remigio, K. G. Dyall, E. Eliav, E. Fasshauer, T. Fleig, L. Halbert, E. D. Hedegård, B. Helmich-Paris, M. Iliaš, C. R. Jacob, S. Knecht, J. K. Laerdahl, M. L. Vidal, M. K. Nayak, M. Olejniczak, J. M. H. Olsen, M. Pernpointner, B. Senjean, A. Shee, A. Sunaga, and J. N. P. van Stralen, The Journal of Chemical Physics 152, (2020).
- Dyall and Faegri (2007) K. G. Dyall and K. Faegri, Introduction to Relativistic Quantum Chemistry (Oxford University Press, 2007).
- Knecht et al. (2022) S. Knecht, M. Repisky, H. J. Aa. Jensen, and T. Saue, The Journal of Chemical Physics 157, (2022).
- Dyall (2023) K. G. Dyall, “Dyall dz, tz, and qz basis sets for relativistic electronic structure calculations,” (2023).
@@
- Yuan, Visscher, and Gomes (2022) X. Yuan, L. Visscher, and A. S. P. Gomes, The Journal of Chemical Physics 156, (2022).
@@
- mrc (2024) “MRCC,” (2024), MRCC, a quantum chemical program suite written by M. Kállay, P. R. Nagy, D. Mester, L. Gyevi-Nagy, J. Csóka, P. B. Szabó, Z. Rolik, G. Samu, J. Csontos, B. Hégely, Á. Ganyecz, I. Ladjánszki, L. Szegedy, B. Ladóczki, K. Petrov, M. Farkas, P. D. Mezei, and R. A. Horváth. See www.mrcc.hu.
@@
- Shee, Visscher, and Saue (2016) A. Shee, L. Visscher, and T. Saue, The Journal of Chemical Physics 145, (2016).
- Gomes and Lyakh (2023) A. Gomes and D. I. Lyakh, “Relmbdev/mb-autogen: mb-autogen,” (2023).
@@
- Usov et al. (2024) D. P. Usov, Y. S. Kozhedub, V. V. Meshkov, A. V. Stolyarov, N. K. Dulaev, A. M. Ryzhkov, I. M. Savelyev, V. M. Shabaev, and I. I. Tupitsyn, Physical Review A 109, (2024).
- Saue et al. (2024) T. Saue et al., “HAMP-vQED,” (2024).
- Repisky et al. (2020) M. Repisky, S. Komorovsky, M. Kadek, L. Konecny, U. Ekström, E. Malkin, M. Kaupp, K. Ruud, O. L. Malkina, and V. G. Malkin, The Journal of Chemical Physics 152, (2020).

### v2

@@ IV.3 Application: CO
The CO molecule was chosen due to the interesting aspects of its triple bond stretching, studied in recent works of the Shabaev groupUsov et al. (2024) and Koput.Koput (2024) In the former work, the influence of quantum electrodynamics (QED) corrections is analyzed which is relevant for the HAMP-vQED projectSaue et al. (2024) for which the present code was created. These are minute effects and their study requires highly accurate methods.
@@
As expected, single-reference methods struggle at large bond distances. The nonparallelity of CCSD and CCSDT becomes apparent at the bond distance of Å (Fig. 10) and the CC has convergence issues from Å on. This is also visible for the rightmost two points for CCSDTQ in Fig. 11.
A deeper study of the behavior of higher-order CC upon dissociation, and especially analyzing the minimum CC excitation level to break the bond successfully is beyond the scope of this manuscript, and will be a subject for our further work, once index-permutation symmetry of tensors is fully implemented.
@@ V CONCLUSIONS
tenpi is an open-source code generator published under BSD-3 license.Brandejs and Saue (2024) tenpi addresses the complex challenge of implementing coupled cluster for modern GPU based HPC platforms in a software-architecture sound way.
The code generated from tenpi is the first working implementation of CCSDT and CCSDTQ in DIRAC. Excellent weak scaling behavior is demonstrated with up to 1200 GPUs. For massively parallel CCSD, the performance is comparable to the hand-tuned code with the measured overhead of only 30%, which is an improvement over the 80% overhead of the existing generated CCSD. The code generation and optimization itself is fast relative to similar toolchains up to quadruple level.
Overall, the tenpi framework provides users with an elegant and systematic way of implementing the coupled cluster methods while hiding under the hood most of the complexity related to the equations, parallelization and intermediates. A simple python interface is featured, while the output is a highly optimized Fortran code that controls the parallel tensor library.
@@
The outlook for tenpi includes the development of CCSD(T)Raghavachari et al. (2013) and CCSDT(Q)Kowalski and Piecuch (2000) for distributed-GPU platforms, and addressing the current limitations like the lack of spatial-symmetry support and integration of DIRAC with tensor software that supports index-permutation symmetry. Inclusion of internode communication and real-world performance data into the cost model for intermediate optimization will be studied. The method development can be boosted by a symbolic interface to Mathematica. tenpi can also serve as a standalone production CC package, if the integral generation is interfaced directly to an existing code, like ReSpect.Repisky et al. (2020)
Besides the above, this work maps the latest developments in relevant tensor software and shows how is the transition to a systematic development pipeline necessary to sustain modern coupled cluster codes.
@@ Acknowledgements.
- Dongarra (2023) J. Dongarra, “A Not So Simple Matter of Software,” https://www.hpcwire.com/2022/11/16/jack-dongarra-a-not-so-simple-matter-of-software/ (2023).
@@
- Woolston (2022) C. Woolston, “Why science needs more research software engineers,” http://dx.doi.org/10.1038/d41586-022-01516-2 (May 31 2022).
- Lyakh (2019) D. I. Lyakh, International Journal of Quantum Chemistry 119, 25926 (2019).
@@
- Papadopoulos (2019) A. Papadopoulos, Necessity of Code Generators in Electronic-Structure Theories and their Infrastructure, Master’s thesis, University of Amsterdam, Amsterdam, Netherlands (2019).
@@
- The Mathworks, Inc. (2024) The Mathworks, Inc., “MATLAB,” https://mathworks.com/products/matlab.html (2024).
- Waterloo Maple, Inc. (2024) Waterloo Maple, Inc., “Maple,” https://www.maplesoft.com/products/Maple/ (2024).
- Wolfram Research, Inc. (2024) Wolfram Research, Inc., “Mathematica, Version 14.0,” https://www.wolfram.com/mathematica (2024), Champaign, IL.
- The Khronos Group, Inc. (2021) The Khronos Group, Inc., “SYCL,” https://www.khronos.org/sycl/ (2021).
- The OpenACC Organization (2022) The OpenACC Organization, “OpenACC,” https://www.openacc.org (2022).
@@
- Brandejs et al. (2024) J. Brandejs, P. Bientinesi, L. Visscher, A. Gomes, and T. Saue, “CECAM Workshop on Tensor Contraction Library Standardization,” https://tensor.sciencesconf.org/ (2024).
- Brandejs and Saue (2024) J. Brandejs and T. Saue, “tenpi: Tensor Programming Interface,” https://gitlab.com/janbrandejs/tenpi (2024).
- Saue et al. (2024) T. Saue et al., “HAMP-vQED,” https://dirac.ups-tlse.fr/hamp-vqed/doku.php?id=public:overview (2024).
@@
- Ozog et al. (2013) D. Ozog, J. R. Hammond, J. Dinan, P. Balaji, S. Shende, and A. Malony, in 2013 42nd International Conference on Parallel Processing (IEEE, 2013) p. 30–39.
@@
- Shiozaki (2017) T. Shiozaki, WIREs Computational Molecular Science 8, 1331 (2017).
@@
- Song, Henderson, and Scuseria (2022) R. Song, T. M. Henderson, and G. E. Scuseria, The Journal of Chemical Physics 156, 104105 (2022).
- Saitow, Kurashige, and Yanai (2013) M. Saitow, Y. Kurashige, and T. Yanai, The Journal of Chemical Physics 139, 044118 (2013).
- Neuscamman, Yanai, and Chan (2009) E. Neuscamman, T. Yanai, and G. K.-L. Chan, The Journal of Chemical Physics 130, 124102 (2009).
- Backhouse (2024) O. Backhouse, “ebcc,” https://github.com/BoothGroup/ebcc (2024).
- Köhn, Richings, and Tew (2008) A. Köhn, G. W. Richings, and D. P. Tew, The Journal of Chemical Physics 129, 201103 (2008).
@@
- Quintero-Monsebaiz and Loos (2023) R. Quintero-Monsebaiz and P.-F. Loos, AIP Advances 13, 085035 (2023).
- Lyakh, Ivanov, and Adamowicz (2004) D. I. Lyakh, V. V. Ivanov, and L. Adamowicz, The Journal of Chemical Physics 122, 024108 (2004).
- Wang, Peyton, and Crawford (2022) Z. Wang, B. G. Peyton, and T. D. Crawford, Journal of Chemical Theory and Computation 18, 5479–5491 (2022).
@@
- Yadav et al. (2024) R. Yadav, M. Bauer, D. Broman, M. Garland, A. Aiken, and F. Kjolstad, “Automatic Tracing in Task-Based Runtime Systems,” https://arxiv.org/abs/2406.18111 (2024), arXiv:2406.18111 [cs.DC] .
@@
- Psarras et al. (2022) C. Psarras, L. Karlsson, J. Li, and P. Bientinesi, “The landscape of software for tensor computations,” https://arxiv.org/abs/2103.13756 (2022), arXiv:2103.13756 [cs.MS] .
@@
- Lyakh et al. (2022) D. I. Lyakh, T. Nguyen, D. Claudino, E. Dumitrescu, and A. J. McCaskey, Frontiers in Applied Mathematics and Statistics 8, 838601 (2022).
- Sharir, Shashua, and Carleo (2022) O. Sharir, A. Shashua, and G. Carleo, Physical Review B 106, 205136 (2022).
@@
- Springer and Bientinesi (2017) P. Springer and P. Bientinesi, “Design of a high-performance GEMM-like Tensor-Tensor Multiplication,” https://arxiv.org/abs/1607.00145 (2017), arXiv:1607.00145 [cs.MS] .
- cuT (2019) “NVIDIA cuTENSOR: A High-Performance CUDA Library For Tensor Primitives,” https://docs.nvidia.com/cuda/cutensor/latest/index.html (2019).
- Bientinesi et al. (2022) P. Bientinesi, D. Ham, F. Huang, P. H. J. Kelly, P. S. Sadayappan, and E. Stow, Dagstuhl Reports , 1 (2022).
@@
- Murray et al. (2023) R. Murray, J. Demmel, M. W. Mahoney, N. B. Erichson, M. Melnichenko, O. A. Malik, L. Grigori, P. Luszczek, M. Dereziński, M. E. Lopes, T. Liang, H. Luo, and J. Dongarra, “Randomized Numerical Linear Algebra : A Perspective on the Field With an Eye to Software,” https://arxiv.org/abs/2302.11474 (2023), arXiv:2302.11474 [math.NA] .
- TAPP working group (2024) TAPP working group, “Tensor Algebra Processing Primitives,” https://github.com/TAPPorg/tensor-interfaces (2024).
@@
- tal (2018) “TAL-SH,” https://github.com/DmitryLyakh/TAL_SH (2018).
- Mutlu et al. (2023) E. Mutlu, A. Panyala, N. Gawande, A. Bagusetty, J. Glabe, J. Kim, K. Kowalski, N. P. Bauman, B. Peng, H. Pathak, J. Brabec, and S. Krishnamoorthy, The Journal of Chemical Physics 159, 024801 (2023).
@@
- Visscher et al. (2024) L. Visscher, H. J. Aa. Jensen, R. Bast, A. S. P. Gomes, T. Saue, I. A. Aucar, V. Bakken, J. Brandejs, C. Chibueze, J. Creutzberg, K. G. Dyall, S. Dubillard, U. Ekström, E. Eliav, T. Enevoldsen, E. Faßhauer, T. Fleig, O. Fossgaard, L. Halbert, E. D. Hedegård, T. Helgaker, B. Helmich-Paris, J. Henriksson, M. van Horn, M. Iliaš, C. R. Jacob, S. Knecht, S. Komorovský, O. Kullie, J. K. Lærdahl, C. V. Larsen, Y. S. Lee, N. H. List, H. S. Nataraj, M. K. Nayak, P. Norman, A. Nyvang, M. Olejniczak, J. Olsen, J. M. H. Olsen., A. Papadopoulos, Y. C. Park, J. K. Pedersen, M. Pernpointner, J. V. Pototschnig, R. Di Remigio Eikås, M. Repiský, K. Ruud, P. Salek, B. Schimmelpfennig, B. Senjean, A. Shee, J. Sikkema, A. Sunaga, J. Thyssen, J. van Stralen, M. L. Vidal, S. Villaume, O. Visser, T. Winther, S. Yamamoto, and X. Yuan, “DIRAC24,” https://zenodo.org/doi/10.5281/zenodo.10680560 (2024).
- Saue et al. (2020) T. Saue, R. Bast, A. S. P. Gomes, H. J. Aa. Jensen, L. Visscher, I. A. Aucar, R. Di Remigio, K. G. Dyall, E. Eliav, E. Fasshauer, T. Fleig, L. Halbert, E. D. Hedegård, B. Helmich-Paris, M. Iliaš, C. R. Jacob, S. Knecht, J. K. Laerdahl, M. L. Vidal, M. K. Nayak, M. Olejniczak, J. M. H. Olsen, M. Pernpointner, B. Senjean, A. Shee, A. Sunaga, and J. N. P. van Stralen, The Journal of Chemical Physics 152, 204104 (2020).
- NEC (2024) NEC, “NEC receives order for next-generation supercomputer system from Japan’s National Institutes for Quantum Science and Technology and National Institute for Fusion Science,” https://www.nec.com/en/press/202411/global_20241113_02.html (2024).
- Hewlett Packard Enterprise (2023) (HPE) Hewlett Packard Enterprise (HPE), “Cray MPICH,” https://www.mpich.org/static/docs/slides/2023-sc-bof/HPE.pdf (2023).
- Sucher (1980) J. Sucher, Phys. Rev. A 22, 348 (1980).
- Almoukhalalati et al. (2016) A. Almoukhalalati, S. Knecht, H. J. A. Jensen, K. G. Dyall, and T. Saue, The Journal of Chemical Physics 145, 074104 (2016).
- Jensen (2005) H. J. Aa. Jensen, (2005), Douglas–Kroll the Easy Way, Talk at Conference on Relativistic Effects in Heavy Elements - REHE, Mülheim, Germany, April, 2005. Available at https://doi.org/10.6084/m9.figshare.12046158.
- Kutzelnigg and Liu (2005) W. Kutzelnigg and W. Liu, J. Chem. Phys. 123, 241102 (2005).
- Iliaš and Saue (2007) M. Iliaš and T. Saue, J. Chem. Phys. 126, 064102 (2007).
- Liu and Peng (2009) W. Liu and D. Peng, J. Chem. Phys. 131, 031104 (2009).
- Hess et al. (1996) B. A. Hess, C. M. Marian, U. Wahlgren, and O. Gropen, Chem. Phys. Lett. 251, 365 (1996).
- (116) B. Schimmelpfennig, “AMFI, an atomic mean-field spin-orbit integral program,” University of Stockholm, Stockholm, Sweden; 1996.
- Knecht et al. (2022) S. Knecht, M. Repisky, H. J. Aa. Jensen, and T. Saue, The Journal of Chemical Physics 157, 114106 (2022).
- Dyall (2023) K. G. Dyall, “Dyall dz, tz, and qz basis sets for relativistic electronic structure calculations,” https://doi.org/10.5281/zenodo.7574629 (2023).
@@
- Yuan, Visscher, and Gomes (2022) X. Yuan, L. Visscher, and A. S. P. Gomes, The Journal of Chemical Physics 156, 224108 (2022).
@@
- mrc (2024) “MRCC,” https://www.mrcc.hu/ (2024), MRCC, a quantum chemical program suite written by M. Kállay, P. R. Nagy, D. Mester, L. Gyevi-Nagy, J. Csóka, P. B. Szabó, Z. Rolik, G. Samu, J. Csontos, B. Hégely, Á. Ganyecz, I. Ladjánszki, L. Szegedy, B. Ladóczki, K. Petrov, M. Farkas, P. D. Mezei, and R. A. Horváth. See www.mrcc.hu.
@@
- Gomes and Lyakh (2023) A. Gomes and D. I. Lyakh, “Relmbdev/mb-autogen: mb-autogen,” https://zenodo.org/record/8094430 (2023).
@@
- Usov et al. (2024) D. P. Usov, Y. S. Kozhedub, V. V. Meshkov, A. V. Stolyarov, N. K. Dulaev, A. M. Ryzhkov, I. M. Savelyev, V. M. Shabaev, and I. I. Tupitsyn, Physical Review A 109, 042822 (2024).
- Koput (2024) J. Koput, Journal of Chemical Theory and Computation 20, 9041–9047 (2024).
- Raghavachari et al. (2013) K. Raghavachari, G. W. Trucks, J. A. Pople, and M. Head-Gordon, Chemical Physics Letters 589, 37–40 (2013).
- Kowalski and Piecuch (2000) K. Kowalski and P. Piecuch, The Journal of Chemical Physics 113, 18–35 (2000).
- Repisky et al. (2020) M. Repisky, S. Komorovsky, M. Kadek, L. Konecny, U. Ekström, E. Malkin, M. Kaupp, K. Ruud, O. L. Malkina, and V. G. Malkin, The Journal of Chemical Physics 152, 184101 (2020).